OpenStudy (anonymous):

DG's Plumbing and Heating charges $50 plus $65 per hour for emergency service. Bill remembers being billed just over $250 for an emergency call. How long to the nearest hour was the plumber at Bill's house?

OpenStudy (anonymous):

Can anyone help

OpenStudy (anonymous):

I need help with the steps to solve this

OpenStudy (espex):

Looking at this, can you guess where to start?

OpenStudy (anonymous):

No sir

OpenStudy (espex):

What are the two elements that determine how much I would pay to have this plumbing company do work for me?

OpenStudy (anonymous):

the plumbing and heating charges

OpenStudy (espex):

Of course the charges, but what specifically about the charges? The problem statement gives you this info, "DG's Plumbing and Heating charges $50 plus $65 per hour for emergency service." What are the two components to the charge?

OpenStudy (anonymous):

the house call and the emergency call?

OpenStudy (espex):

Yes, but what are the "numbers" associated with that, remember that this is a math problem.

OpenStudy (anonymous):

60 and 55

OpenStudy (espex):

Yes, exactly.

OpenStudy (espex):

Okay, so what does it cost JUST to have them show up, to make a house call?

OpenStudy (anonymous):

50 dollars

OpenStudy (espex):

Yes, nice work. Okay, and what if they show up AND do 1 hour of work?

OpenStudy (anonymous):

would it be 115

OpenStudy (espex):

Perfect, now, how did you get that number? Write it out mathematically.

OpenStudy (anonymous):

I added both the 50 and 65

OpenStudy (espex):

Good, so how about mathematically, how would you write that out?

OpenStudy (anonymous):

50x+65

OpenStudy (espex):

Where did the x come from?

OpenStudy (espex):

What are you multiplying?

OpenStudy (anonymous):

aw this isn't an equation

OpenStudy (espex):

Technically it is an expression since there is no =

OpenStudy (anonymous):

aw ok

OpenStudy (espex):

So try it again, how would you represent the process you used mathematically?

OpenStudy (anonymous):

can you give me an example of an expression

OpenStudy (espex):

All of these are expressions. 1+2 5-2 12*2 4/1

OpenStudy (anonymous):

ok I thought it was something different so 50+65

OpenStudy (espex):

Excellent, that is exactly right. So now, how much would it cost if I had the plumber show up AND do 2 hours of work?

OpenStudy (anonymous):

180

OpenStudy (espex):

Good, can you write out, mathematically, how you got it?

OpenStudy (anonymous):

50+130=180

OpenStudy (espex):

Good, how could you break up the hourly charge?

OpenStudy (anonymous):

50+1 hour(65)=110 50+2 hours(130)=180

OpenStudy (espex):

That is a good start, what are you expressing with the 2 hours(130)?

OpenStudy (anonymous):

the emergency service

OpenStudy (espex):

What about writing it like this: 50+1*(65)=110, 50+2*(65)=180? You are expressing that there is a $50 charge plus $65 for each hour.

OpenStudy (anonymous):

Aw ok

OpenStudy (espex):

How about if I had them come over AND do 5 hours of work?

OpenStudy (anonymous):

would I multiply by 5

OpenStudy (espex):

Yes, like you did with the 1 and 2

OpenStudy (anonymous):

so it would be 250 and 325

OpenStudy (espex):

Where did you get the 250?

OpenStudy (anonymous):

I multiplied 50*5

OpenStudy (espex):

Why? Did you do that when you had these equations? 50+1*(65)=110, 50+2*(65)=180

OpenStudy (anonymous):

no

OpenStudy (espex):

Okay, so why would you do it now? :)

OpenStudy (anonymous):

I wouldn't multiply I would add

OpenStudy (espex):

So what would your equation look like?

OpenStudy (anonymous):

so would I add 50 dollars 5 times and 65 dollars 5 times?

OpenStudy (espex):

You have "times" in there twice, what components do you have to pay for each of the 5 hours?

OpenStudy (anonymous):

50 and 65

OpenStudy (espex):

What you are doing here is looking at what you have to pay for each hour that the service is there. Just to have them come out costs you $50, then, every hour you are charged $65. You are not charged a call-out fee every hour, you only pay that once. That is why you "add" the call-out fee.

OpenStudy (espex):

Think about what is actually happening, and try to apply that to your equation to see if it fits.

OpenStudy (anonymous):

aw ok so it's only 65

OpenStudy (espex):

Yes, because that is the hourly rate, that is what you are charged for each hour they are there. You do not need to pay a call-out every hour.

OpenStudy (espex):

What would the three equations look like if I asked you to tell me what it cost to have the plumbers there for 2, 3, and 4 hours?

OpenStudy (espex):

I believe that you need to work on the basics a little more. This site has games and tips from pre-algebra up through calculus. Give it a look: http://www.coolmath.com/index.html